# Quasiperiodic Quadrupole Insulators

### v3

@@ Abstract
Higher-order topological insulators are an intriguing new family of topological states that host lower-dimensional boundary states. Concurrently, quasiperiodic systems have garnered significant interest due to their complex localization and topological properties. In this work we study the impact of quasiperiodic modulations on the paradigmatic Benalcazar-Bernevig-Hughes model, which hosts topological insulating phases with zero-energy corner modes. We find that the topological properties are not only robust to the quasiperiodic modulation, but can even be enriched. In particular, we unveil the first instance of a quasiperiodic induced second-order topological insulating phase. Furthermore, in contrast with disorder, we find that quasiperiodic modulations can induce multiple reentrant topological transitions, showing an intricate sequence of localization properties. Our results open a promising avenue for exploring the rich interplay between higher-order topology and quasiperiodicity.
@@ I Introduction
This paper is organized as follows: in Sec. II we introduce the quasiperiodic Benalcazar-Bernevig-Hughes model and present the methods used to numerically study the diverse phases that emerge under QP modulations; in Sec. III we display and discuss our results in a organized fashion; in Sec. IV we discuss and reinforce our main conclusions; in Appendix A we show calculations of the density of states (DOS) obtained via kernel polynomial methods (KPM) for each relevant observed phases and in Appendix. B we perform the same analysis done throughout this work but for a constant QP modulation strength ($W$) cut (varying the inter unit cell hopping $\gamma$); in Appendix C some remarks are made regarding phase twists; and in Appendix. D we expand on the results of phase shifts dependence in the QPQI phase.
@@ II.0.1 Model
where $\beta$ is an irrational number. This choice of QP modulation ensures that chiral symmetry remains preserved for any QP modulation strength ($W$). Furthermore, this model also preserves time reversal and charge conjugation symmetries. For a generic phase shift $\bm{\phi}=(\phi_{x},\phi_{y})$, the QP modulation breaks the mirror and $C_{4}$ crystalline symmetries. Therefore, this symmetry is broken for the finite-size systems that we study. Nonetheless, it should be recovered in the thermodynamic limit, when there should be no $\bm{\phi}$ dependence of the bulk properties.
@@
In finite systems the introduction of phase shifts $\boldsymbol{\phi}=\left(\phi_{x},\phi_{y}\right)$ and phase twists $\boldsymbol{\theta}=(\theta_{x},\theta_{y})$ can alter the energy spectra of the system. The dependence of the energy spectra on $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$ has a direct correlation with the localization properties of the corresponding eigenfunctions, which remain unchanged by different choices of $(\boldsymbol{\phi},\boldsymbol{\theta})$ , as discussed in Ref. (Gonçalves et al., 2023a). In fact, taking simultaneous averages over $(\boldsymbol{\phi},\boldsymbol{\theta})$ in IPR and IPRk calculations, culminates in smoother results with better scalings with system size $N=L_{x}L_{y}$. Regarding topological and spectral properties the same argument does not hold, and in finite systems a spectral gap can appear for particular choices of a boundary twist. In this manner, for spectral gap dependent results [^2] These results include direct calculations of the spectral gap and topological properties we fixed a twist of 00 or $\pi$ depending if the system size is even or odd respectively (see Appendix C). As for dependences in phase shifts, ballistic regimes will have no dependence on the choice of phase shifts, and thus we average our results over phase shifts sampled from a random uniform distribution ($\phi_{x},\phi_{y}\in[0,2\pi[$).
@@
where $\psi_{\alpha}(\boldsymbol{R})=c_{\mathbf{R},\alpha}^{\dagger}\left|0\right\rangle$. As the system approaches the thermodynamic limit, any bulk dependence on phase twists and shifts should vanish .
In the clean limit ($W=0$) the system preserves translational invariance and we re-obtain the BBH model (Benalcazar et al., 2017a) , described by the following Bloch Hamiltonian:
@@
For $\left|\gamma\right|<1$($\left|\gamma\right|>1$) the system is in a topological (trivial) phase defined by $q_{xy}=0.5$ ($q_{xy}=0$). As for the topological phase, the system displays fourfold zero energy corner modes that give rise to fractional corner charges $Q_{i}=\pm 0.5$. Since this quadrupole insulating phase is a second order topological insulator (SOTI), the boundary is a first order topological insulator. In fact, from the boundary perspective the corner modes appear as localized edge states. In this manner, we can characterize the topological phase by calculating the boundary polarizations with the nested Wilson loop approach introduced and discussed in Refs. (Benalcazar et al., 2017a, b). It was shown that in the presence of quantizing crystalline symmetries ($C_{4}$, and reflection along $x$ and $y$, $M_{x}$ and $M_{y}$), the quadrupole insulating phase obeys $q_{xy}=p_{x}=p_{y}=0.5$, and that anti-commuting mirror symmetries are needed to have a Wannier gap. For any given finite value of $W$ these crystalline symmetries are broken for a finite system with generic $\bm{\phi}$, however, it has recently been shown that the quadrupole moment is equally quantized by chiral symmetry (Yang et al., 2021; Li et al., 2020) and that a new set of $\mathbb{Z}$ topological invariants arise, known as the multipole chiral numbers (MCNs) (Benalcazar and Cerjan, 2022). Although this new characterization is shown for systems falling in the AIII symmetry class, it should be valid for any of the chiral symmetric classes (AIII, BDI, CII), since BDI and CII classes display more symmetry than the AIII class. These chiral symmetric quadrupole phases are topological and display the same properties as their crystalline counterparts, such has zero energy corner modes, fractional corner charges, quantized boundary polarizations and bulk quadrupole moments.
@@ III.4 Opening the Boundaries
Properties of QPQI phases - Until now a proper characterization of the QPQI phases has not been provided. In the paragraphs that follow, we aim to fully define these quasiperiodic induced regime. We start with their definition: a QPQI phase is a topological quadrupole insulating phase that is not adiabatically connected to any of the clean limit phases while displaying a quasiperiodic induced bulk-spectral gap (or is gapless but displays localized states at the Fermi level, as discussed in Appendix B).
@@ IV Conclusions
We have studied the phase diagram of a quasiperiodic quadrupole chiral insulator. The higher-order topological phases of the parent model were found to be stable upon the addition of quasiperiodicity. More interestingly, quasiperiodicity was also found to induce multiple reentrant topological transitions into quadrupole insulating phases, expanding the known class of chiral symmetric quadrupole insulators introduced in Ref. (Khalaf et al., 2021; Benalcazar and Cerjan, 2022)
Similarly to disorder-induced transitions into higher-order topological Anderson insulators (Li et al., 2020; Yang et al., 2021; Lóio et al., 2024), quasiperiodic-induced topological transitions occurring under the usual gap closing and reopening mechanism were observed. However, in contrast to the disordered case, the eigenstates in the gap edge were found to be ballistic across the transition, as in the homogeneous limit. Also in contrast with the disordered case, multiple quasiperiodic-induced topological transitions were found.
A further interesting question that deserves further attention concerns the interplay between the quasiperiodicity and higher-order topology. Quasiperiodicity alone can induce a rich edge physics, which was not the focus of our study. However, these edge states can have a non-trivial interplay with the zero energy corner modes that deserves further exploration.
@@ Acknowledgements
Appendix A Edge and Corner States via DOS
Appendix B Constant $W$ Cut

### v4

@@ Abstract
Higher-order topological insulators are an intriguing new family of topological states that host lower-dimensional boundary states. Concurrently, quasiperiodic systems have garnered significant interest due to their complex localization and topological properties. In this work we study the impact of chiral symmetry preserving quasiperiodic modulations on the paradigmatic Benalcazar-Bernevig-Hughes model, which hosts topological insulating phases with zero-energy sublattice-polarized modes. We find that the topological properties are not only robust to the quasiperiodic modulation, but can even be enriched. In particular, we unveil the first instance of a quasiperiodic induced second-order topological insulating phase. Furthermore, in contrast with disorder, we find that quasiperiodic modulations can induce multiple reentrant topological transitions, showing an intricate sequence of localization properties. Our results open a promising avenue for exploring the rich interplay between higher-order topology and quasiperiodicity.
@@ I Introduction
This paper is organized as follows: in Sec. II we introduce the quasiperiodic Benalcazar-Bernevig-Hughes model and present the methods used to numerically study the diverse phases that emerge under QP modulations; in Sec. III we display and discuss our results in a organized fashion; in Sec. IV we discuss and reinforce our main conclusions; in Appendix B we show calculations of the density of states (DOS) obtained via kernel polynomial methods (KPM) for each relevant observed phases and in Appendix. A we perform the same analysis done throughout this work but for a constant QP modulation strength ($W$) cut (varying the inter unit cell hopping $\gamma$); in Appendix C some remarks are made regarding phase twists; and in Appendix. D we expand on the results of phase shifts dependence in the QPQI phase.
@@ II.0.1 Model
where $\beta$ is an irrational number. This choice of QP modulation ensures that chiral symmetry remains preserved for any QP modulation strength ($W$). Furthermore, this model also preserves time reversal and charge conjugation symmetries. In periodic finite systems, a generic phase shift $\bm{\phi}=(\phi_{x}\mod{2\pi},\phi_{y}\mod{2\pi})$, breaks the mirror and $C_{4}$ crystalline symmetries. However, for a set of phase shifts $\phi_{i}\in\left\{\pi\beta(2m_{i}+1)\mod{2\pi}|m_{i}\in\mathbb{Z}\right\}$ (for $L$ even) or $\phi_{i}\in\left\{2\pi\beta m_{i}\mod{2\pi}|m_{i}\in\mathbb{Z}\right\}$ (for $L$ odd) the potential is shifted by $\mathbf{R}=m_{x}\hat{\mathbf{e}}_{x}+m_{y}\hat{\mathbf{e}}_{y}$, and the original crystalline symmetries are preserved. . Nonetheless, in the thermodynamic limit the mirror, inversion, and $C_{4}$ symmetries are recovered for a generic $\bm{\phi}=(\phi_{x},\phi_{y})$. In contrast, finite open systems only retain said symmetries around the center of the lattice. In a finite open quasiperiodic system, the crystalline symmetries are only compatible with the potential for a single choice of $\boldsymbol{\phi}=\left(\phi_{x},\phi_{y}\right)$, which aligns the geometric centers of the lattice and the potential  [^2] For a lattice with geometric center at $(m_{x},m_{y})$, symmetry is retained for $\left(\phi_{x},\phi_{y}\right)=(\pi\beta(2m_{x}+1),\pi\beta(2m_{y}+1))\mod{2\pi}$ (for even $L$), $\left(\phi_{x},\phi_{y}\right)=(2\pi\beta m_{x},2\pi\beta m_{y})\mod{2\pi}$ (for odd $L$)..
The BBH model can be interpreted as having fractional corner charges protected by mirror symmetries and no chiral symmetry or sublattice-polarized zero modes protected by chiral symmetry. In this work, we choose the latter interpretation since chiral symmetry is the only preserved symmetry for a general choice of potential parameters and open or periodic boundaries.
@@
In finite systems the introduction of phase shifts $\boldsymbol{\phi}=\left(\phi_{x},\phi_{y}\right)$ and phase twists $\boldsymbol{\theta}=(\theta_{x},\theta_{y})$ can alter the energy spectra of the system. The dependence of the energy spectra on $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$ has a direct correlation with the localization properties of the corresponding eigenfunctions, which remain unchanged by different choices of $(\boldsymbol{\phi},\boldsymbol{\theta})$, as discussed in Ref. (Gonçalves et al., 2023a). In fact, taking simultaneous averages over $(\boldsymbol{\phi},\boldsymbol{\theta})$ in IPR and IPRk calculations, culminates in smoother results with better scalings with system size $N=L_{x}L_{y}$. Regarding topological and spectral properties the same argument does not hold, and in finite systems a spectral gap can appear for particular choices of a boundary twist. In this manner, for spectral gap dependent results [^3] These results include direct calculations of the spectral gap and topological properties we fixed a twist of 00 or $\pi$ depending if the system size is even en or odd respectively (see Appendix C). In the thermodynamic limit, the $\boldsymbol{\phi}$-dependence vanishes for the bulk properties. For a quasiperiodic potential in a square lattice, the phase shifts appear as twists in the reciprocal space Hamiltonian, thus phase shifts are transformed to boundary conditions in reciprocal space and have no impact on the thermodynamic limit. More generally, in an infinite quasiperiodic system, all local configurations of the quasiperiodic potential are realized, and introducing a phase shift leads to no change in the bulk states. In finite open systems, localized and critical regimes depend on the choice of phase shifts, which set the local configuration of the quasiperiodic potential. Moreover, edge properties, such as the edge spectrum, retain a dependence on phase shifts even in the thermodynamic limit since different shifts change the local configurations at the boundaries, no matter the size of the system. A simple way of probing all the possible configurations in the infinite system is to average over different phase shifts. However, this strategy in finite open systems should never preserve the symmetry on average. The averaging mechanism in quasiperiodic systems is comparable to that performed in disordered systems, where average symmetries of disordered ensembles can be similar to the exact ones (Chaou et al., 2024). It is plausible that the same holds in quasiperiodic systems, even though the unique interplay between quasiperiodic potentials and edge physics may raise some concerns.
@@
where $\psi_{\alpha}(\boldsymbol{R})=c_{\mathbf{R},\alpha}^{\dagger}\left|0\right\rangle$. In the clean limit ($W=0$) the system preserves translational invariance and we re-obtain the BBH model (Benalcazar et al., 2017a) , described by the following Bloch Hamiltonian:
@@
For $\left|\gamma\right|<1$($\left|\gamma\right|>1$) the system is in a topological (trivial) phase defined by $q_{xy}=0.5$ ($q_{xy}=0$). As for the topological phase, the system displays fourfold zero energy corner modes that give rise to fractional corner charges $Q_{i}=\pm 0.5$. Since this quadrupole insulating phase is a second order topological insulator (SOTI), the boundary is a first order topological insulator. In fact, from the boundary perspective the corner modes appear as localized edge states. In this manner, we can characterize the topological phase by calculating the boundary polarizations with the nested Wilson loop approach introduced and discussed in Refs. (Benalcazar et al., 2017a, b). It was shown that in the presence of quantizing crystalline symmetries ($C_{4}$, and reflection along $x$ and $y$, $M_{x}$ and $M_{y}$), the quadrupole insulating phase obeys $q_{xy}=p_{x}=p_{y}=0.5$, and that anti-commuting mirror symmetries are needed to have a Wannier gap. For any given finite value of $W$ these crystalline symmetries are broken for a finite system with generic $\bm{\phi}$, however, it has recently been shown that the quadrupole moment is equally quantized by chiral symmetry (Yang et al., 2021; Li et al., 2020) and that a new set of $\mathbb{Z}$ topological invariants arise, known as the multipole chiral numbers (MCNs) (Benalcazar and Cerjan, 2022). Although this new characterization is shown for systems falling in the AIII symmetry class, it should be valid for any of the chiral symmetric classes (AIII, BDI, CII), since BDI and CII classes display more symmetry than the AIII class. Chiral symmetric quadrupole phases are higher order topological phases and can exhibit zero energy corner modes, quantized edge polarizations and bulk quadrupole moments.
@@ III.4 Opening the Boundaries
Properties of QPQI phases - Until now a proper characterization of the QPQI phases has not been provided. In the paragraphs that follow, we aim to fully define these quasiperiodic induced regime. We start with their definition: a QPQI phase is a topological quadrupole insulating phase that is not adiabatically connected to any of the clean limit phases while displaying a quasiperiodic induced bulk-spectral gap (or is gapless but displays localized states at the Fermi level, as discussed in Appendix A).
@@ IV Conclusions
We studied the phase diagram of a quasiperiodic quadrupole chiral insulator. The higher-order topological phases of the parent model were found to be stable upon the addition of quasiperiodicity. More interestingly, quasiperiodicity induces multiple reentrant topological transitions into quadrupole insulating phases, expanding the known class of chiral symmetric quadrupole insulators introduced in Ref. (Khalaf et al., 2021; Benalcazar and Cerjan, 2022)
Similarly to disorder-induced transitions into higher-order topological Anderson insulators (Li et al., 2020; Yang et al., 2021; Lóio et al., 2024), quasiperiodic-induced topological transitions occur under the usual bulk gap closing mechanism were observed compatible an intrinsic HOTI. However, in contrast to the disordered case, the eigenstates in the gap edge were found to be ballistic across the transition, as in the homogeneous limit, and multiple quasiperiodic-induced topological transitions were found.
The main result of this work is the identification of a reentrant topological phase in which corner states coexist with gapless edge states, showcasing the rich interplay between quasiperiodicity and intrinsic higher-order topology. Lifting the degeneracy of the corner modes leads to the emergence of quantized fractional corner charges, characteristic of a canonical quadrupole moment. The gapless edge states can interact nontrivially with the zero-energy corner modes, suggesting a complex edge–corner interplay that warrants further investigation.
@@ Acknowledgements
Appendix A Constant $W$ Cut